A 70–kg dairy heifer is placed in an indirect metabolism chamber and the following results are obtained:

Oxygen consumption: 373 L/d

Carbon dioxide produced: 330 L/d

Urine N excreted: 9 g N/d

Given the following standards:

Standard Values for Protein

g of Urine N x 6.25 = grams of protein

1 g protein: 0.96 L O2 consumed

1 g protein: 0.77 L CO2 produced

1 g protein: 4.3 kcal heat produced

Non-Protein RQ (CO2/O2)

O2 (kcal/L)

% Heat from CHO

% Heat from Fat

0.70

4.686

0

100

0.80

4.801

33.4

66.6

0.90

4.924

67.5

32.5

How much heat was generated in kcal per day?

The correct answer is 1813.
Answer #1

grams of protein= 9g of Urine N x 6.25 = 56.25 g

O2 Consumed by protein = 56.25 x 0.96L= 54 L/d

CO2 Produced= 56.25 x 0.77 L= 43.31 L/d

Heat generated from protein = 56.25 x 4.3 kcal = 241.87 kcal.

O2 consumed by non-protein substance = 373 - 54 = 319 L/d

CO2 Produced by non-protein substance = 330 - 43.31 = 286.69 L/d

Non-protein RQ = 286.69 / 319 = 0.90

So, heat produced from non-protein substance = 319 x 4.924 = 1570.756 kcal

Total heat generated per day = 241.87 + 1570.756 = 1813 kcal